### v2

@@ Abstract
Visual Grounding (VG) tasks, such as referring expression detection and segmentation tasks are important for linking visual entities to context, especially in complex reasoning tasks that require detailed query interpretation. This paper explores VG beyond basic perception, highlighting challenges for methods that require reasoning like human cognition. Recent advances in large language methods (LLMs) and Vision-Language methods (VLMs) have improved abilities for visual comprehension, contextual understanding, and reasoning. These methods are mainly split into end-to-end and compositional methods, with the latter offering more flexibility. Compositional approaches that integrate LLMs and foundation models show promising performance but still struggle with complex reasoning with language-based logical representations. To address these limitations, we propose NAVER, a compositional visual grounding method that integrates explicit probabilistic logic reasoning within a finite-state automaton, equipped with a self-correcting mechanism. This design improves robustness and interpretability in inference through explicit logic reasoning. Our results show that NAVER achieves SoTA performance comparing to recent end-to-end and compositional baselines. The code is available at https://github.com/ControlNet/NAVER.
@@ 3.3 Logic Generation ($\boldsymbol{S_{L}}$)
To translates the extracted visual information and the text query into explicit logic representations for reasoning, the Logic Generation state is designed to generate the ProbLog [14] logic code for reasoning. ProbLog is selected as the logic representation because it supports probabilistic logic, allowing the system to reason about the uncertainties that naturally arise in visual tasks (e.g., overlapping entities or ambiguous attributes).
@@ 4.3 Quantitative Analysis
Compositional Code-bison [43] 44.4 38.2 - -
@@
Performance Analysis. The previous compositional baselines [44, 21] compute the grounding performance excluding the runtime errors caused by LLM’s uncontrollable output. The scores reported in the benchmarks are not reliable in real applications due to the high failure rate. We conduct a fair comparison with previous compositional baselines [21, 44] by using the same LLM (GPT-4o Mini [35]), grounding model (GLIP-Large [26], and SAM [22]) in the segmentation task. We report the IoU scores including/excluding the runtime errors and the runtime failure rate in Table 5. The original ViperGPT lacks the validation in inference, leading to a high failure rate. HYDRA reduces the error rate by introducing incremental reasoning and a controller to validate the plan, but still reports a 10.3% failure rate. Our proposed method NAVER produces a more reliable pipeline (0.3% runtime failure rate) with the validation for the full-chain inference, where each state verifies the intermediate outputs. We find the main failure reason is due to the ChatGPT content policy which misclassifies some queries such as “man cut off” as violent text and fails to respond, however this query asks for a man only half of whose body appears in the image. Furthermore, comparing the performance ignoring the runtime error, we still observe a noticeable improvement of NAVER comparing to the baselines with the same foundation models.
Complexity Analysis. Efficiency results for NAVER and compositional baselines are provided in Table 6, which includes average runtime, token usage, and cost per sample. For a fair comparison, we use the same foundation models and LLM for the experiments. From the results, we observe NAVER is more efficient than HYDRA [21] with higher performance and lower runtime failure rate. Although ViperGPT [44] has lower time complexity, this advantage is offset by reduced performance due to its lack of validation. These results demonstrate that NAVER’s flexible automaton design reduces unnecessary computations by dynamically adapting the pipeline.
@@
Time # LLM Tokens LLM Cost
Method (Second) Input Output (USD)
ViperGPT [44] 2.02 4169 29 0.00064
HYDRA [21] 14.93 19701 618 0.00332
NAVER (Ours) 5.96 2083 39 0.00034
@@ 5 Conclusion
This supplementary material provides additional details about the proposed method NAVER. In particular, it provides details about the additional quantitative comparison result with both accuracy and IoU in section 6, the ablation studies for VLM in section 7, the analysis of self-correction mechanisms in section 8, the performance analysis for query length in section 9, and the prompts used for the LLMs and VLMs in section 10.
@@ 6 Additional Quantitative Comparison
In this section, we report both the acurracy and Intersection over Union (IoU) performance for the referring expression detection task. Table 10 presents a quantitative comparison on the RefCOCO, RefCOCO+, RefCOCOg [20], and Ref-Adv [1] datasets. Additionally, we include results for the compositional methods using GroundingDINO as the VFM, marked with ‡. We observed that for each configuration, our proposed method NAVER outperforms the other compositional methods using the same foundation models and also outperform the grounding methods themselves. These results further validate that our approach consistently achieves superior IoU performance over existing baselines.
@@ 9 Query Length Analysis
We analyze the impact of text query length on performance, comparing NAVER with baselines across different query lengths. The results, shown in Figure 4, indicate that NAVER consistently reaches SoTA performance comparing to all baselines regardless of query length. While most baseline methods experience a performance decline as query length increases, the extent of decrease varies. Notably, older models like GLIP [26] show a significant performance drop for longer queries, suggesting difficulties in handling complex text inputs. In contrast, NAVER maintains stable performance, showing its robustness in processing longer and more complex queries.
@@ 10 LLM and VLM Prompts
NAVER employs LLMs and VLMs in five different roles as described in section 3: as a caption generator VLM in the Perception state, an entity extractor LLM in the Perception state, a logic query generator LLM in the Logic Generation state, a relation recognizer VLM in the Logic Generation state, and an answerer in the Answering state. The prompts utilized for each role are carefully crafted to generate precise outputs aligned with the requirements of each state, ensuring accurate visual grounding and robust reasoning. Prompts 10.1 to 10.5 show the detailed prompt templates for each role.
  • Prompt 10.1: This prompt instructs the VLM to generate descriptive captions $I_{c}$ for the input image $I$. The generated captions represent the visual content in textual form, which serves as input for the entity extractor LLM in subsequent processing.
  • Prompt 10.2: This prompt is utilized in the Perception state to guide the LLM in identifying entity categories $C$ relevant to the query $Q$ and the caption $I_{c}$. The output is a list of entity categories, which forms the basis for visual reasoning in the subsequent states.
  • Prompt 10.3: This prompt guides the logic query generator LLM to convert the textual query $Q$ into a ProbLog logic query in the Logic Generation state. The output logic query is then utilized in the ProbLog interpreter to perform probabilistic logic reasoning.
  • Prompt 10.4: This prompt guides the relation recognizer VLM to identify relations $R$ between entities $E$ in the Logic Generation state. These recognized relations are used to construct logic expressions, as the foundation for probabilistic logic reasoning.
  • Prompt 10.5: This prompt is used in the Answering state to guide the answerer VLM in validating whether the identified target $Y_{L1}$ satisfies the conditions of the query $Q$. It requests a binary response (“Yes” or “No”) to confirm whether the top-ranked candidate from the Logic Reasoning state fulfills all query requirements, ensuring only valid results are returned as final outputs.
Prompt 10.1: Captioner VLM ⟨image⟩Please describe the image in detail.
Prompt 10.2: Entity Extractor LLM You’re an AI assistant designed to find detailed information from image.
@@
Prompt 10.3: Logic Query Generator LLM You’re an AI assistant designed to generate the ProbLog code (a logic programming language similar to Prolog).
@@
Prompt 10.4: Relation Recognizer VLM ⟨image⟩You’re an AI assistant designed to find the relations of objects in the given image.
@@
Prompt 10.5: Answerer VLM ⟨image⟩You’re an image analyst designed to check if the highlighted object in the image meets the query description.

### v3

@@ Abstract
Visual Grounding (VG) tasks, such as referring expression detection and segmentation tasks are important for linking visual entities to context, especially in complex reasoning tasks that require detailed query interpretation. This paper explores VG beyond basic perception, highlighting challenges for methods that require reasoning like human cognition. Recent advances in large language methods (LLMs) and Vision-Language methods (VLMs) have improved abilities for visual comprehension, contextual understanding, and reasoning. These methods are mainly split into end-to-end and compositional methods, with the latter offering more flexibility. Compositional approaches that integrate LLMs and foundation models show promising performance but still struggle with complex reasoning with language-based logical representations. To address these limitations, we propose NAVER, a compositional visual grounding method that integrates explicit probabilistic logic reasoning within a finite-state automaton, equipped with a self-correcting mechanism. This design improves robustness and interpretability in inference through explicit logic reasoning. Our results show that NAVER achieves SoTA performance compared to recent end-to-end and compositional baselines. The code is available at https://github.com/ControlNet/NAVER.
@@ 3.3 Logic Generation ($\boldsymbol{S_{L}}$)
To translate the extracted visual information and the text query into explicit logic representations for reasoning, the Logic Generation state is designed to generate the ProbLog [15] logic code for reasoning. ProbLog is selected as the logic representation because it supports probabilistic logic, allowing the system to reason about the uncertainties that naturally arise in visual tasks (e.g., overlapping entities or ambiguous attributes).
@@ 4.3 Quantitative Analysis
GPT-4o-2024-05-13 [36] 30.5 26.2 - -
Qwen2.5-VL-72B [4] 94.6 92.2 90.3 -
Compositional
Code-bison [44] 44.4 38.2 - -
@@
Performance Analysis. The previous compositional baselines [45, 22] compute the grounding performance excluding the runtime errors caused by LLM’s uncontrollable output. The scores reported in the benchmarks are not reliable in real applications due to the high failure rate. We conduct a fair comparison with previous compositional baselines [22, 45] by using the same LLM (GPT-4o Mini [36]), grounding model (GLIP-Large [27], and SAM [23]) in the segmentation task. We report the IoU scores including/excluding the runtime errors and the runtime failure rate in Table 5. The original ViperGPT lacks the validation in inference, leading to a high failure rate. HYDRA reduces the error rate by introducing incremental reasoning and a controller to validate the plan, but still reports a 10.3% failure rate. Our proposed method NAVER produces a more reliable pipeline (0.3% runtime failure rate) with the validation for the full-chain inference, where each state verifies the intermediate outputs. We find the main failure reason is due to the ChatGPT content policy which misclassifies some queries such as “man cut off” as violent text and fails to respond, however this query asks for a man only half of whose body appears in the image. Furthermore, comparing the performance ignoring the runtime error, we still observe a noticeable improvement of NAVER compared to the baselines with the same foundation models.
@@ 5 Conclusion
Acknowledgments
This research is sponsored by the DARPA Assured Neuro Symbolic Learning and Reasoning (ANSR) program under award number FA8750-23-2-1016.
@@ Acknowledgments
This supplementary material provides additional details about the proposed method NAVER. In particular, it provides details about the additional quantitative comparison result with both accuracy and IoU in section 6, the complexity analysis in section 7, the ablation studies for VLM in section 8, the analysis of self-correction mechanisms in section 9, the performance analysis for query length in section 10, the ablation studies for captioner in section 11, and the prompts used for the LLMs and VLMs in section 12.
@@ 6 Additional Quantitative Comparison
In this section, we report both the accuracy and Intersection over Union (IoU) performance for the referring expression detection task. Table 9 presents a quantitative comparison on the RefCOCO, RefCOCO+, RefCOCOg [21], and Ref-Adv [1] datasets. Additionally, we include results for the compositional methods using GroundingDINO as the VFM, marked with ‡. We observed that for each configuration, our proposed method NAVER outperforms the other compositional methods using the same foundation models and also outperform the grounding methods themselves. These results further validate that our approach consistently achieves superior IoU performance over existing baselines.
7 Complexity Analysis
Efficiency results for NAVER and compositional baselines are provided in Table 10, which includes average runtime, token usage, and cost per sample. For a fair comparison, we use the same foundation models and LLM for the experiments. From the results, we observe NAVER is more efficient than HYDRA [22] with higher performance and lower runtime failure rate. Although ViperGPT [45] has lower time complexity, this advantage is offset by reduced performance due to its lack of validation. These results demonstrate that NAVER’s flexible automaton design reduces unnecessary computations by dynamically adapting the pipeline.
Time # LLM Tokens LLM Cost
Method (Seconds) Input Output (USD)
ViperGPT [45] 2.02 4169 29 0.00064
HYDRA [22] 14.93 19701 618 0.00332
NAVER (Ours) 5.96 2083 39 0.00034
@@ 8 Ablation Studies for VLMs
Captioner ECE RefCOCO RefCOCO+ RefCOCOg RefAdv
VLM 91.6 87.5 80.2 65.6
VLM LLM 96.2 92.8 91.6 75.4
Table 12: Ablations for the captioner and type of entity category extractor (ECE). All results are accuracy for referring expression detection task.
@@ 10 Query Length Analysis
We analyze the impact of text query length on performance, comparing NAVER with baselines across different query lengths. The results, shown in Figure 4, indicate that NAVER consistently reaches SoTA performance compared to all baselines regardless of query length. While most baseline methods experience a performance decline as query length increases, the extent of decrease varies. Notably, older models like GLIP [27] show a significant performance drop for longer queries, suggesting difficulties in handling complex text inputs. In contrast, NAVER maintains stable performance, showing its robustness in processing longer and more complex queries.
11 Ablations for Captioner
In the perception state, NAVER first converts the image into a rich caption and then lets an LLM-based entity-category extractor (ECE) decide which object classes are relevant to the query. An intuitive alternative is to skip captioning and ask the VLM to predict categories directly. Table 12 shows that this seemingly simpler choice causes a consistent accuracy drop of 4.6% on RefCOCO, 5.3% on RefCOCO+, 11.4% on RefCOCOg, and 9.8% on Ref-Adv. The larger gap on the more complex datasets (RefCOCOg, Ref-Adv) suggests that a textual scene summary helps the downstream logic generator reason about complex descriptions. The clear improvement brought by the Captioner justifies keeping it in the final system.
@@ 12 LLM and VLM Prompts
NAVER employs LLMs and VLMs in five different roles as described in section 3: as a caption generator VLM in the Perception state, an entity extractor LLM in the Perception state, a logic query generator LLM in the Logic Generation state, a relation recognizer VLM in the Logic Generation state, and an answerer in the Answering state. The prompts utilized for each role are carefully crafted to generate precise outputs aligned with the requirements of each state, ensuring accurate visual grounding and robust reasoning. Prompts 12.1 to 12.5 show the detailed prompt templates for each role.
  • Prompt 12.1: This prompt instructs the VLM to generate descriptive captions $I_{c}$ for the input image $I$. The generated captions represent the visual content in textual form, which serves as input for the entity extractor LLM in subsequent processing.
  • Prompt 12.2: This prompt is utilized in the Perception state to guide the LLM in identifying entity categories $C$ relevant to the query $Q$ and the caption $I_{c}$. The output is a list of entity categories, which forms the basis for visual reasoning in the subsequent states.
  • Prompt 12.3: This prompt guides the logic query generator LLM to convert the textual query $Q$ into a ProbLog logic query in the Logic Generation state. The output logic query is then utilized in the ProbLog interpreter to perform probabilistic logic reasoning.
  • Prompt 12.4: This prompt guides the relation recognizer VLM to identify relations $R$ between entities $E$ in the Logic Generation state. These recognized relations are used to construct logic expressions, as the foundation for probabilistic logic reasoning.
  • Prompt 12.5: This prompt is used in the Answering state to guide the answerer VLM in validating whether the identified target $Y_{L1}$ satisfies the conditions of the query $Q$. It requests a binary response (“Yes” or “No”) to confirm whether the top-ranked candidate from the Logic Reasoning state fulfills all query requirements, ensuring only valid results are returned as final outputs.
Prompt 12.1: Captioner VLM ⟨image⟩Please describe the image in detail.
Prompt 12.2: Entity Extractor LLM You’re an AI assistant designed to find detailed information from image.
@@
Prompt 12.3: Logic Query Generator LLM You’re an AI assistant designed to generate the ProbLog code (a logic programming language similar to Prolog).
@@
Prompt 12.4: Relation Recognizer VLM ⟨image⟩You’re an AI assistant designed to find the relations of objects in the given image.
@@
Prompt 12.5: Answerer VLM ⟨image⟩You’re an image analyst designed to check if the highlighted object in the image meets the query description.